OpenStudy (mollynicole2015):

Manoj and Tamar are arguing about how a number trick they heard goes. Tamar tells their friend Andrew to think of a number, multiply it by five, and subtract three from the result. Then Manoj tells Andrew to think of a number, add five, and multiply the result by three. Andrew says that whichever way he does the trick he gets the same answer. What was Andrew's number? Justify your answer.

OpenStudy (jackthegreatest):

let x = Andrew's number let y = the result 5x - 3 = y (x + 5) 3 = y 3x + 15 = y 3x + 15 = 5x - 3 18 = 2x x = 9